OpenStudy (anonymous):

21. Choose the better or best buy. 2 pt for $1.59 8 oz for $0.39 6 oz for $0.30 22. Choose the better or best buy. 5 lb for $0.99 1 lb for $0.75 25 lb for $4.89

OpenStudy (shane_b):

For 21 note that there 16 ounces in a pint so write out each ratio. Each result will be the price per unit...so the smaller price per unit is the better the value is. \[\frac{$1.59}{16*2}=?\]\[\frac{$0.30}{6}=?\]\[\frac{$0.39}{8}=?\]For 22 it's basically the same thing.